# To be or not to be, but where?

### v1

@@ 6.2 You can’t always get what you want
where $\kappa=\sqrt{32\pi G}$. This theory is diffeomorphic invariant We will consider a fixed flat background with perturbations on top, so we expand $g_{\mu\nu}=\eta_{\mu\nu}+\kappa h_{\mu\nu}$, where $h_{\mu\nu}$ is a canonically normalized field. Under infinitesimal linearized diffeomorphisms, the fields transform as
@@
Using (6.2) and (6.4) it is straightforward to show the invariance of $\Phi$ at leading order in $\kappa$. As we will discuss more in Sec. (7), we expect that any final theory of quantum gravity will have an effective perturbative regime that falls into the theory being studied here, and we argue that much of the current physical phenomenology is modeled, tacitly or explicitly, under this linear perturbative regime.
@@
Now that we have obtained dressed gravitational operators that are invariant under linear diffs, do they satisfy microcausality? Unfortunately, not. As we had before, the commutator of $\Phi$ between two spacelike regions depends on $[V(x),V(y)]$ as in (5.8), but inside of $V^{\mu}$ there is the metric and its derivatives, including its time-derivative. What was crucial for the QED example was that the dressing was made with the gauge field, which is the same object that was quantized, but that is not the case for gravity. Typically, in canonical quantization of gravity, both non-perturbative and linearized gravity, we quantize the metric instead, which its necessity can be traced back to the introduction of the Ricci scalar (both the Riemann tensor and Ricci tensor can be fully written only in terms of the connection) [^22] Here, we see a very clear motivation to consider Loop Quantum Gravity where the connection, expressed in terms of Ashtekar variables, is quantized instead of the metric. For a recent discussion, [54, and references therein].. You might wonder about the other type of dressed observable (5.9), but those are not available due to the absence of negatively charged particles under gravity.
The evidence against the existence of gauge-invariant local observables in quantum gravity is compelling. Already three decades ago, [55] demonstrated that in closed universes, invariant observables cannot be constructed as local functions of the canonical variables, linking the absence of local observables to topological symmetries. More recently, [56] argues that while classical gravitational observables can be microcausal, this property is limited at quantum scales, with significant implications for black hole information. Additionally, [41] shows that any operator with nonzero Poincaré charges must be gravitationally dressed, implying nonlocality and the failure of defining local subsystems in gravity and [53] highlights the conflict of defining mutually independent subsystems via standard tensor products or operator algebras in gravitational contexts. For more references, see [57, section Observables and references therein] and [58].
@@ 7.1 Early Universe Cosmology
We do not have direct access to the very early universe, but among all the different cosmological observations that offer us a window into its early stages, the Cosmic Microwave Background (CMB) radiation takes the main role. Spread out throughout the whole universe, it provides an aged snapshot of the Universe approximately 380 000 years after the Big Bang. It is a snapshot in the sense that this background radiation was produced during a short phase where the light in the early Universe went from propagating in an opaque medium, where the photons were constantly interacting with electrons and baryons, to a transparent one where they could freely propagate until reaching our current detectors. It is aged because the Universe has evolved since this transition, and the detected radiation distribution today carries memories of the Universe’s history when compared to its expected profile back when it was emitted. For example, one of the memories impinges upon the whole radiation spectrum, which has been redshifting since its emission, and today is detected at a temperature of $2.7$ K as opposed to its initial temperature of around $3000$ K.
@@
As we discussed in Sec 3.3, if we are talking about temperature it must mean that we are close to thermal equilibrium. That is the case for the CMB, which presents an overall spectrum very close to that of a black body, with small temperature fluctuations of order $10^{-5}$ K. These fluctuations are also part of the spectrum’s memory, as they reflect small density variations in the early Universe. One of the initial surprising features of the CMB was exactly this proximity to thermal equilibrium. The medium transition from opaque to transparent happens as the local temperature decreases enough such that, on average, photons cannot excite electrons out of proton’s electric potential anymore, allowing them to form stable hydrogen. This local temperature depends on the amount of local energy densities. Observing thermal equilibrium in the CMB indicates that the energy density was very evenly distributed both homogeneously and isotropically, otherwise different parts of the Universe would reach this transition at different times, leading to larger variations of the temperature of the detected photons depending on which part of the Universe we are probing (or, literally, looking at). This even distribution of energy density is particularly puzzling because, under the standard Big Bang evolution, the CMB corresponds to causally disconnected patches of the universe. Therefore, we need to explain how the Universe’s energy distribution was so even at the time the CMB was emitted [^23] For more quantitative details, see [64, 65]..
@@
Moreover, even if we could argue that the initial state of the Universe corresponded to perfect homogeneity due to unknown properties of quantum gravity, we would still need to explain why the fluctuations in the CMB are not random, but instead have correlations over very large distances that were not causally connected, so the light did not have enough time to travel that much since the beginning of the Big Bang. In fact, without assuming any new physics between the Big Bang and the emission of the CMB, we can show that points separated by more than two degrees on the sky seem to not have been in causal contact.
@@
Thus, currently, the early universe paradigm is based on apparent violations of macrocausality due to a high degree of homogeneity of the radiation in the CMB together with seemingly super-causal correlations. Meanwhile, the current paradigm offers a solution based on inflation, which is typically modeled by a scalar field sourcing the fast-expanding cosmological background, while combined quantization of the field and spacetime fluctuations are ultimately responsible for the primordial density fluctuations. Hence, the paradigm falls into LEQG.
@@
This can potentially introduce some internal inconsistency. The reason is that the very reason to introduce new physics in the early universe is based on assumptions (macrocausality) that are not a given within the background theoretical framework (LEQG).
@@
Based on this argument, a future direction is to study the early universe using non-local gauge-invariant observables dressed by gravitational gauge fields. It is known that dressing a field theory creates a new Fock space from which scattering amplitudes can be analyzed, in particular the states are defined over a new vacuum. Thus, basically we have either a set of observables with the undressed vacuum, or a dressed set of observables with the dressed vacuum. This is reminiscent of the setup of the Unruh effect, or Hawking radiation, where different observers, one inertial and one accelerating, have access to different sets of observables and vacua, and the vacuum state of one is seen as a thermal bath to the other. Could our observations of the CMB seen from the local perspective [^24] Since gravity is not modelled in the detection, it is as if we are framing the observations from the point of view of the undressed theory. look thermal exactly because of the mismatch between these vacua? [^25] A similar argument can be made in the context of the cosmological constant problem, as the calculations making the problem explicit typically rely on the Fock space defined by the free quantum field theory [66]. It would be interesting to check whether the zero-point energies of the dressed modes contribute differently than the ones defined by the free theory..
7.2 Gravitationally-Induced-Entanglement (GIE) Experiments
@@
The quantum nature of gravity remains a topic of debate [67, 68, and references therein]. Thus, it seems more than justified the hype around the possibility of probing whether gravity is non-classical in the lab [60, 61]. Originally, the experimental setup involves two masses, each in a spatial superposition, interacting only through gravity. Due to their spatial superposition, each pair of trajectories induces a different unitary phase due to the interaction Hamiltonian, which depends on the relative distance of the trajectories, different for all the pairs. Thus, if the initial state was a product state between the superimposed spatial states of each mass, after their interaction each term will have a different phase, and the overall state can no longer be decomposed as a product state. Following [61], if the initial state is considered
@@
The experiment described above is known as the bipartite model, which is typically considered a Newtonian approximation of the tripartite model. For this model, gravity appears as a mediator subsystem, such that the masses interact only directly with it. Thus, for instance, instead of $|L\rangle_{1}|L\rangle_{2}$, we have $|L\rangle_{1}|\gamma_{LL}\rangle|L\rangle_{2}$ which evolves to $e^{i\phi_{LL}}|L\rangle_{1}|\gamma_{LL}\rangle|L\rangle_{2}$. As each of the $|\gamma\rangle$ states represents a different classical gravitational configuration, there is a superposition of classical gravitational states. What is crucial to the tripartite model assuming interactions between the masses only mediated by the subsystem $|\gamma\rangle$ is that, if the $|\gamma\rangle$ is a classical subsystem (the set of its observables define a commuting subalgebra), then one can show that no entanglement can be produced [69, see for an overview]. Hence, were the experiment performed, witnessing the entanglement between the two masses would be strong evidence of the quantum nature of gravity.
@@
At least that is the common lore. Critical analyses of the modeling and implication of a positive detection have already been discussed [70]. In [63], it has been argued that such experiments are probing the quantum nature of gravity similarly as the quantized cosmological perturbations are being probed by the CMB observations under the inflationary paradigm discussed above. They highlight two key differences: i) cosmological perturbations focus on field fluctuations, whereas GIE experiments involve cat states, which are significantly more quantum (measured by non-classicality such as Wigner negativity), and ii) GIE experiments explore a new, untested regime of physics, namely, the gravitation of delocalized masses.
@@
Here we want to call attention to yet two other aspects underlying the modeling of such experiments that, as far as I know, have not been pointed out so far. Due to the low energy regime and reliance on perturbative quantum gravity, these experiments fall within the LEQG framework. However, as discussed in Sec 2, a well-defined tensor product structure, necessary for composing different systems, is only ensured in the presence of microcausality. Moreover, the mass states above are also defined as excitations of some free matter field [61, see supplementary material], which given our discussion in Sec. 6.2, are not gauge invariant under linear diffeormorphisms. Thus, the consistent use of LEQG to describe such experiments, believed to be fundamental to any possible conclusion to be derived from either positive or negative entanglement detections, seems to require an improvement of their description both in terms of the factorization understanding of the Hilbert space as well as the type of states that are considered [^26] There is yet another issue that has to do with state preparation. In the context of field theories, the idea behind a product state is generalized to what is called statistical independence [31, 3], which guarantees that any pair of state preparations made by two different experimenters in two spacelike regions can be seen as a preparation of the system as a whole. Achieving statistical independence requires both microcausality and the split property, which together ensure a well-defined tensor product structure and thus product states. Once again, LEQG might seem to violate microcausality and it remains unclear whether quantum gravity would satisfy anything like the split property [71, 72, 73]..
@@ 8 Single-world unitary quantum mechanics
Our discussion about the GIE experiments invites us towards a much broader shift in our conceptual perspective. The breakdown of microcausality in LEQG, and the subsequent loss of a well-defined tensor product structure, undermines the concept of quantum-gravitational subsystems, which due to the universal nature of gravity applies to everything. This shift may offer a new perspective on the measurement problem, suggesting that quantum subsystems are not invariant under measurements or generic Hamiltonian evolution [^27] This idea was first explored in [74] by modeling an apparatus and a quantum system, both as qubits. Although the mathematical implementation was limited, some of the conceptual framework was developed there, though not in the context of LEQG and microcausality violations. More recently, thanks to Jan Glowacki and Tein van der Lugt, I became aware of [75], where similar ideas are pursued over solid mathematical foundations. Perhaps, the discussion presented here offers a possible answer to their final question concerning a unified story between their framework and quantum gravity.. Just as the relativization of time and space was crucial for the development of special relativity, the relativization of quantum subsystems might be essential for resolving the measurement problem and advancing towards a final theory of quantum gravity.
@@ 8.1 Quantum subsystems are not time-invariant
Crucially, note that from the quartet $(\mathcal{H},\rho_{0},B(\mathcal{H}),H,)$, now we have factorized each of these elements, and associated with the factorization $\mathcal{T}$ are the coefficients,
$\{p_{n},a_{0},a_{I},b_{J},c_{IJ}\}=\{p_{n}(\mathcal{T}),a_{0}(\mathcal{T}),a_{I}(\mathcal{T}),b_{J}(\mathcal{T}),c_{IJ}(\mathcal{T})\}\,,$ (8.9)
which parametrize the states and the Hamiltonian, thus kinematics and dynamics of the theory, in a given TPS. Their dependence on the TPS can be simply understood by the fact that our starting point did not have any of these coefficients present, thus they are contingent on the TPS chosen. On the other hand, they are not all independent from one another, as they are all connected to the class of unitary map $\mathcal{T}$ being considered, and its unitary nature will impose constraints on these parameters.
@@
  • For $d/D\lll 1$ (which is absolutely the case), there are many inequivalent TPSs $\mathcal{T}_{i}$. In fact, as they are connected by global unitaries that can typically be continuously parametrized, we will assume any two TPS are related by continuous deformations and parametrized by $\tau$: $\mathcal{T}(\tau)$ [^30] Despite introducing this infinity to parameterize all the different TPS that fix the dimensionality of the Hilbert spaces of the apparatus and EE, it pales in comparison with all the possible different TPS that can be generically considered, e.g., bipartite with the same dimensionality for each subsystem or non-bipartite ones. Another way of understanding this is by taking the Hamiltonian perspective from [5], where one always considers a fixed TPS but then varies the Hamiltonian, akin to the coefficients in (8.9) evolving.;
@@
We will show that by putting together these assumptions we can arrive at the following conclusion:
@@
Let’s start by showing this and then we discuss the assumptions and conclusion in light of a close analogy to finding local inertial frames in general relativity. Whenever $\mathcal{T}(\tau_{0})\in\{\mathcal{T}_{M}\}$, we have a product state between what we call apparatus and everything else, so we can write the state as
@@
and $\bar{\mathcal{O}}^{\rm app^{\prime}}_{I}=U^{-1}\tilde{\mathcal{O}}^{\rm app}_{I}U$. Thus, the local state is one of the rotated apparatus’ eigenstates with eigenvalue $o_{I,n}^{\rm app^{\prime}}$. Crucially, we do not know which one. Then, it is easy to show that
@@
Therefore, by identifying the appropriate TPS and locally rotating/aligning the apparatus, we see that local measurement outcomes are actually the global invariants that we started with. Crucially, to arrive at this point, besides our assumptions (to be scrutinized), we only relied on global and local unitary transformations. We expect that the probabilistic nature of the theory arises from our lack of knowledge about the specific set of unitaries required to reach this point, such that probabilities will be epistemic. This will be closely investigated in the future. In any case, given that all considerations involved only unitary transformations (both for TPS updates and time evolution) and we still arrive at a description of the apparatus with single outcomes related to the global invariants, we term this yet-to-be-finalized theory single-world unitary quantum mechanics. Under $\mathcal{T}_{M}$, local measurements can be associated with single outcomes, even though all transformations considered were unitary.

### v2

@@ 6.2 You can’t always get what you want
where $\kappa=\sqrt{32\pi G}$. This theory is diffeomorphic invariant. We will consider a fixed flat background with perturbations on top, so we expand $g_{\mu\nu}=\eta_{\mu\nu}+\kappa h_{\mu\nu}$, where $h_{\mu\nu}$ is a canonically normalized field. Under infinitesimal linearized diffeomorphisms, the fields transform as
@@
Using (6.2) and (6.4), it is straightforward to show the invariance of $\Phi$ at leading order in $\kappa$. As we will discuss more in Sec. (7), we expect that any final theory of quantum gravity will have an effective perturbative regime that falls into the theory being studied here, and we argue that much of the current physical phenomenology is modeled, tacitly or explicitly, under this linear perturbative regime.
@@
Now that we have obtained dressed gravitational operators that are invariant under linear diffs, do they satisfy microcausality? Unfortunately, not. As we had before, the commutator of $\Phi$ between two spacelike regions depends on $[V(x),V(y)]$ as in (5.8), but inside of $V^{\mu}$ there is the metric and its derivatives, including its time-derivative. What was crucial for the QED example was that the dressing was made with the gauge field, which is the same object that was quantized, but that is not the case for gravity. Typically, in the covariant quantization of gravity, both non-perturbative and linearized gravity, we quantize the metric instead, which its necessity can be traced back to the introduction of the Ricci scalar (both the Riemann tensor and Ricci tensor can be fully written only in terms of the connection) [^25] Here, we see a very clear motivation to consider Loop Quantum Gravity where the connection, expressed in terms of Ashtekar variables, is quantized instead of the metric. For a recent discussion, [60, and references therein].. You might wonder about the other type of dressed observable (5.9), but those are not available due to the absence of negatively charged particles under gravity.
The evidence against the existence of gauge-invariant local observables in quantum gravity is compelling. Already three decades ago, [61] demonstrated that in closed universes, invariant observables cannot be constructed as local functions of the canonical variables, linking the absence of local observables to topological symmetries. More recently, [62] argues that while classical gravitational observables can be microcausal, this property is limited at quantum scales, with significant implications for black hole information. Additionally, [47] shows that any operator with nonzero Poincaré charges must be gravitationally dressed, implying nonlocality and the failure of defining local subsystems in gravity and [59] highlights the conflict of defining mutually independent subsystems via standard tensor products or operator algebras in gravitational contexts. For more references, see [63, section Observables and references therein] and [64].
@@ 7.1 Early Universe Cosmology
We do not have direct access to the very early universe, but among all the different cosmological observations that offer us a window into its early stages, the Cosmic Microwave Background (CMB) radiation takes the main role. Spread out throughout the whole universe, it provides an aged snapshot of the Universe approximately 380,000 years after the Big Bang. It is a snapshot in the sense that this background radiation was produced during a short phase where the light in the early Universe went from propagating in an opaque medium, where the photons were constantly interacting with electrons and baryons, to a transparent one where they could freely propagate until reaching our current detectors. It is aged because the Universe has evolved since this transition, and the detected radiation distribution today carries memories of the Universe’s history when compared to its expected profile back when it was emitted. For example, one of the memories impinges upon the whole radiation spectrum, which has been redshifting since its emission, and today is detected at a temperature of $2.7$ K as opposed to its initial temperature of around $3000$ K.
@@
As we discussed in Sec 3.3, if we are talking about temperature, it must mean that we are close to thermal equilibrium. That is the case for the CMB, which presents an overall spectrum very close to that of a black body, with small temperature fluctuations of order $10^{-5}$ K. These fluctuations are also part of the spectrum’s memory, as they reflect small density variations in the early Universe. One of the initial surprising features of the CMB was exactly this proximity to thermal equilibrium. The medium transition from opaque to transparent happens as the local temperature decreases enough such that, on average, photons cannot excite electrons out of the proton’s electric potential anymore, allowing them to form stable hydrogen. This local temperature depends on the amount of local energy density. Observing thermal equilibrium in the CMB indicates that the energy density was very evenly distributed both homogeneously and isotropically; otherwise, different parts of the Universe would reach this transition at different times, leading to larger variations of the temperature of the detected photons depending on which part of the Universe we are probing (or, literally, looking at). This even distribution of energy density is particularly puzzling because, under the standard Big Bang evolution, the CMB corresponds to causally disconnected patches of the universe. Therefore, we need to explain how the Universe’s energy distribution was so even at the time the CMB was emitted [^26] For more quantitative details, see [70, 71]..
@@
Moreover, even if we could argue that the initial state of the Universe corresponded to perfect homogeneity due to unknown properties of quantum gravity, we would still need to explain why the fluctuations in the CMB are not random, but instead have correlations over very large distances that were not causally connected, so the light did not have enough time to travel that much since the beginning of the Big Bang. In fact, without assuming any new physics between the Big Bang and the emission of the CMB, we can show that points separated by more than two degrees on the sky do not seem to have been in causal contact.
@@
Thus, currently, the early universe paradigm is based on apparent violations of macrocausality due to a high degree of homogeneity of the radiation in the CMB together with seemingly super-causal correlations. Meanwhile, the current paradigm offers a solution based on inflation, which is typically modeled by a scalar field sourcing the fast-expanding cosmological background, while the combined quantization of the field and spacetime fluctuations is ultimately responsible for the primordial density fluctuations. Hence, the paradigm falls into LEQG.
@@
This can potentially introduce some internal inconsistency. The reason is that the very reason to introduce new physics in the early universe is based on assumptions (macrocausality) that are not a given within the background theoretical framework (LEQG), as we discussed in Sec. 6.2.
@@
Based on this argument, a future direction is to study the early universe using non-local gauge-invariant observables dressed by gravitational gauge fields. It is known that dressing a field theory creates a new Fock space from which scattering amplitudes can be analyzed; in particular, the states are defined over a new vacuum. Thus, basically we have either a set of observables with the undressed vacuum or a dressed set of observables with the dressed vacuum. This is reminiscent of the setup of the Unruh effect, or Hawking radiation, where different observers, one inertial and one accelerating, have access to different sets of observables and vacua, and the vacuum state of one is seen as a thermal bath to the other. Could our observations of the CMB seen from the local perspective [^27] Since gravity is not modelled in the detection, it is as if we are framing the observations from the point of view of the undressed theory. look thermal exactly because of the mismatch between these vacua? [^28] A similar argument can be made in the context of the cosmological constant problem, as the calculations making the problem explicit typically rely on the Fock space defined by the free quantum field theory [72]. It would be interesting to check whether the zero-point energies of the dressed modes contribute differently than the ones defined by the free theory.
7.2 Gravitationally Induced Entanglement (GIE) Experiments
@@
The quantum nature of gravity remains a topic of debate [73, 74, and references therein]. Thus, it seems more than justified the hype around the possibility of probing whether gravity is non-classical in the lab [66, 67]. Originally, the experimental setup involved two masses, each in a spatial superposition, interacting only through gravity. Due to their spatial superposition, each pair of trajectories induces a different unitary phase due to the interaction Hamiltonian, which depends on the relative distance of the trajectories, different for all the pairs. Thus, if the initial state was a product state between the superimposed spatial states of each mass, after their interaction, each term will have a different phase, and the overall state can no longer be decomposed as a product state. Following [67], if the initial state is considered
@@
The experiment described above is known as the bipartite model, which is typically considered a Newtonian approximation of the tripartite model. For this model, gravity appears as a mediator subsystem, such that the masses interact only directly with it. Thus, for instance, instead of $|L\rangle_{1}|L\rangle_{2}$, we have $|L\rangle_{1}|\gamma_{LL}\rangle|L\rangle_{2}$ which evolves to $e^{i\phi_{LL}}|L\rangle_{1}|\gamma_{LL}\rangle|L\rangle_{2}$. As each of the $|\gamma\rangle$ states represents a different classical gravitational configuration, there is a superposition of classical gravitational states. What is crucial to the tripartite model assuming interactions between the masses only mediated by the subsystem $|\gamma\rangle$ is that, if the $|\gamma\rangle$ is a classical subsystem (the set of its observables define a commuting subalgebra), then one can show that no entanglement can be produced [75, see for an overview] [^29] Previous arguments relied on LOCC-style reasoning [68], typically assuming finite-dimensional systems. Only recently has this gap been closed by showing that any classical mediator, modeled as a commutative unital $C^{*}$-algebra, cannot generate entanglement, even for infinite-dimensional quantum systems and general $C^{*}$-tensor products [76].. Hence, were the experiment performed, witnessing the entanglement between the two masses would be strong evidence of the quantum nature of gravity.
@@
At least that is the common lore. Critical analyses of the modeling and implications of a positive detection have already been discussed [77]. In [69], it has been argued that such experiments are probing the quantum nature of gravity similarly to the quantized cosmological perturbations being probed by the CMB observations under the inflationary paradigm discussed above. They highlight two key differences: i) cosmological perturbations focus on field fluctuations, whereas GIE experiments involve cat states, which are significantly more quantum (measured by non-classicality such as Wigner negativity), and ii) GIE experiments explore a new, untested regime of physics, namely, the gravitation of delocalized masses.
@@
Here, we want to call attention to yet two other aspects underlying the modeling of such experiments that have not been pointed out so far. Due to the low-energy regime and reliance on perturbative quantum gravity, these experiments fall within the LEQG framework. However, as discussed in Sec 2, a well-defined tensor product structure, necessary for composing different systems, is only ensured in the presence of microcausality. Moreover, the mass states above are also defined as excitations of some free matter field [67, see supplementary material], which, given our discussion in Sec. 6.2, are not gauge invariant under linear diffeormorphisms. Thus, the consistent use of LEQG to describe such experiments, believed to be fundamental to any possible conclusion to be derived from either positive or negative entanglement detections, seems to require an improvement of their description both in terms of the factorization understanding of the Hilbert space as well as the type of states that are considered [^30] There is yet another issue that has to do with state preparation. In the context of field theories, the idea behind a product state is generalized to what is called statistical independence [36, 4], which guarantees that any pair of state preparations made by two different experimenters in two spacelike regions can be seen as a preparation of the system as a whole. Achieving statistical independence requires both microcausality and the split property, which together ensure a well-defined tensor product structure and thus product states. Once again, LEQG might seem to violate microcausality, and it remains unclear whether quantum gravity would satisfy anything like the split property [78, 79, 80]..
@@ 8 Single-world unitary quantum mechanics
Our discussion about the GIE experiments invites us towards a much broader shift in our conceptual perspective. The breakdown of microcausality in LEQG, and the subsequent loss of a well-defined tensor product structure, undermines the concept of quantum-gravitational subsystems, which, due to the universal nature of gravity, applies to everything. This shift may offer a new perspective on the measurement problem, suggesting that quantum subsystems are not invariant under measurements or generic Hamiltonian evolution [^31] This idea was first explored in [81] by modeling an apparatus and a quantum system, both as qubits. Although the mathematical implementation was limited, some of the conceptual framework was developed there, though not in the context of LEQG and microcausality violations. More recently, thanks to Jan Glowacki and Tein van der Lugt, I became aware of [82], where similar ideas are pursued over solid mathematical foundations. Another paper that explores similar ideas can be found in [83].. Just as the relativization of time and space was crucial for the development of special relativity, the relativization of quantum subsystems might be essential for resolving the measurement problem and advancing towards a final theory of quantum gravity.
To some extent, this shift should not come as a surprise. A central lesson of general relativity is background independence: the idea that no structure should be fixed and exempt from dynamics. But a fixed tensor product structure in Hilbert space [^32] Admittedly, our discussion so far has overlooked a crucial subtlety, particularly evident in the Heisenberg picture, where time evolution under the Hamiltonian naturally generates a continuous family of inequivalent TPSs [84, 11]. In what follows, we will work with an initially fixed, kinematically selected TPS defined at an initial reference time [5]. Still, a more robust and general approach treats subsystems not in terms of Hilbert space factorization, but rather through the algebraic structure of operator subalgebras and their interrelations, an idea that finds its most precise expression in the formalism of descriptors [85, 86]., or a fixed assignment of algebras to spacetime regions, violates this very principle. Once gravity is subject to quantum principles, it seems natural that such structures must also become dynamical. In fact, just as different spacetime foliations in special relativity are induced kinematically by different inertial observers, the kinematical relativization of subsystems in quantum theory was already emphasized in [87]. What remains open is how to promote this relativization to a dynamical feature of the theory. Here, we aim to outline what such a step might involve.
@@ 8.1 Quantum subsystems are not time-invariant
Crucially, note that the quartet $(\mathcal{H},B(\mathcal{H}),H,\rho_{0})$ is now factorized, and associated with the factorization $\mathcal{T}$ are the coefficients,
$\{p_{n},a_{0},a_{I},b_{J},c_{IJ}\}=\{p_{n}(\mathcal{T}),a_{0}(\mathcal{T}),a_{I}(\mathcal{T}),b_{J}(\mathcal{T}),c_{IJ}(\mathcal{T})\}\,.$ (8.9)
which parametrize the states and the Hamiltonian, thus the kinematics and dynamics of the theory, in a given TPS. Their dependence on the TPS can be simply understood by the fact that our starting point did not have any of these coefficients present, thus they are contingent on the TPS chosen. On the other hand, they are not all independent from one another, as they are all connected to the class of unitary maps $\mathcal{T}$ being considered, and its unitary nature will impose constraints on these parameters.
@@
  • For $d/D\lll 1$ (which is absolutely the case), there are infinitely many inequivalent TPSs $\mathcal{T}_{i}$ [84]. In fact, as they are connected by global unitaries that can be continuously parametrized, we will assume any two TPS are related by continuous deformations and parametrized by $\tau$: $\mathcal{T}(\tau)$ [^35] Despite introducing this infinity to parameterize all the different TPS that fix the dimensionality of the Hilbert spaces of the apparatus and EE, it pales in comparison with all the possible different TPS that can be generically considered, e.g., bipartite with the same dimensionality for each subsystem or non-bipartite ones. Another way of understanding this is by taking the Hamiltonian perspective from [6], where one always considers a fixed TPS but then varies the Hamiltonian, akin to the coefficients in (8.9) evolving. This is akin to how the renormalization group flow changes the values of the coupling parameters modeling interactions as the energy scale changes in the effective field theory paradigm, now understood as the notion of the relevant quantum subsystems being updated depending on which modes are integrated out.;
@@
We will show that by putting together these assumptions, we can arrive at the following conclusion:
@@
Let’s start by showing this, and then we will discuss the assumptions and conclusion in light of a close analogy to finding local inertial frames in general relativity. Whenever $\mathcal{T}(\tau_{0})\in\{\mathcal{T}_{M}\}$, we have a product state between what we call the apparatus and everything else, so we can write the state as
@@
and $\bar{\mathcal{O}}^{\rm app^{\prime}}_{I}=U^{-1}\tilde{\mathcal{O}}^{\rm app}_{I}U$. Thus, the local state is one of the rotated apparatus’ eigenstates with eigenvalue $o_{I,n}^{\rm app^{\prime}}$. Crucially, we do not know which one. Then, it is easy to show that,
@@
Therefore, by identifying the appropriate TPS and locally rotating/aligning the apparatus, we see that local measurement outcomes are actually the global invariants that we started with, closing the gap between the holistic and atomistic ontologies in quantum mechanics at the empirical level. Crucially, to arrive at this point, besides our assumptions (to be scrutinized), we only relied on global and local unitary transformations. We expect that the probabilistic nature of the theory arises from our lack of knowledge about the specific set of unitaries required to reach this point, such that probabilities will be epistemic. This will be closely investigated in the future. In any case, given that all considerations involved only unitary transformations (both for TPS updates and time evolution) and we still arrive at a description of the apparatus with single outcomes related to the global invariants, we term this yet-to-be-finalized theory single-world unitary quantum mechanics. Under $\mathcal{T}_{M}$, local measurements can be associated with single outcomes, even though all transformations considered were unitary.